OpenStudy (anonymous):

Can someone tell me what 8*8 is?

Mehek (mehek14):

add 8 eight times 8*8 = 8+8+8+8+8+8+8+8 = ?

OpenStudy (anonymous):

Just kidding I'm just trolling

pooja195 (pooja195):

Thats really annoying.

Mehek (mehek14):

u wot m8?

pooja195 (pooja195):

Quit spamming m8

OpenStudy (anonymous):

I'm not spamming

OpenStudy (vera_ewing):

Spam

OpenStudy (anonymous):

spam

pooja195 (pooja195):

Funny how mods do nothing about this.

Mehek (mehek14):

^

pooja195 (pooja195):

We need more mods.

OpenStudy (anonymous):

Look up the definition of spam

OpenStudy (anonymous):

I'm not even spamming bruh

pooja195 (pooja195):

Trolling is against the code of conduct.

OpenStudy (anonymous):

And so is your face

OpenStudy (vera_ewing):

spam: inappropriate or irrelevant messages/posts/comments sent on the internet. This post is spam. Please close it.

OpenStudy (anonymous):

64

OpenStudy (vera_ewing):

You're diverting attention away from people who actually need help.

pooja195 (pooja195):

on't spam! Common forms of spam include: Posting a question multiple times within a short period of time. Posting a question that is off topic for a particular group (i.e. posting a physics question in the Mathematics group). Posting a reply that is not relevant to the topic of discussion within that question. Posting questions that are not, in fact, questions (for example, starting a question in the Mathematics group to wish someone "Happy Birthday"). Sending the same chat message repeatedly. Filing false abuse reports on a user repeatedly.

pooja195 (pooja195):

Im sick of mods doing nothing.

OpenStudy (anonymous):

Abuse reported.

OpenStudy (anonymous):

Well geesh calm yourself. You're taking it too seriously

pooja195 (pooja195):

Go help people.

OpenStudy (vera_ewing):

^

pooja195 (pooja195):

Stop wasting time

OpenStudy (anonymous):

If someone uses * instead of "x or "times", that means someone is familiar with multiplication. TLDR: troll post